# Superconducting qubits at the utility scale: the potential and limitations of modularity

### v2

@@ A.1 Glossary of default RRE outputs
    $t^{\text{decode-delay}}_{tot}$ represents the total decoding delay accumulated from all graph consumption steps, which we must add to the overall fault-tolerant hardware time. We assume there are no decoying delays during subgraph hand-over operations due to $t_{\text{tock}}^{\text{decoding}}\leq 8t_{\text{inter}}d$. We can determine the total decoding delay, given that $t_{\text{tock}}^{\text{decoding}}>t_{\text{tock}}^{\text{quantum}}$ and $t_{\text{tock}}^{\text{decoding}}>8Ct$, using $t^{\text{decode-delay}}_{\text{tot}}=(8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{prep}})(t_{\text{tocks}}^{\text{decoding}}-t_{\text{tock}}^{\text{quantum}})+\lceil\frac{t^{\text{delay}}_{\text{distill}}}{8Ct}\rceil(t_{\text{tocks}}^{\text{decoding}}-8Ct)$.
@@
  • Parameter 46: Fault-tolerant wall-time over a single algorithmic step, $t^{\text{hardware}}_{\text{tot}}$
@@
    $t^{\text{hardware}}_{\text{tot}}$ is our estimated upper-bound for the fault-tolerant wall-time accumulated from all non-simultaneous components running in modules on both legs and interconnects of the architecture of section I.3 over a single algorithmic step, $U$. Algorithmic step should not be confused with $n_{\text{widgets}}$ widgetization steps of $U$. An algorithmic step refers to a situation where the user needs $n_{\text{algo-reps}}$ repetitions of $U$ executed on the FTQC. This wall-time includes all the inter and intra tocks required at the consumption stages, delays needed for the graph preparation and $T$-distillation, subgraph handover time, and the decoding delay. In other words, this is our overall time cost, which can be formulated as $t_{\text{hardware}}^{total}=t^{\text{total}}_{\text{intra}}+t^{\text{total}}_{\text{handover-inter}}+t_{\text{decode-delay}}^{\text{total}}$. See section I.4.6 for details.
  • Parameter 47: Total fault-tolerant algorithm time, $t^{\text{FT}}_{\text{total}}$
@@
    RRE allows users to specify the number of times, $n_{\text{algo-reps}}$, they would like to repeat an exact copy of the complete algorithm $U$ back to back. For these cases, the total time is calculated as $t^{\text{FT}}_{\text{total}}=n_{\text{algo-reps}}t_{\text{hardware}}^{total}$.
  • Parameter 48: Total energy consumption, $E_{\text{tot}}$
@@
    We report estimates for $E_{\text{tot}}$, the upper-bound on the total energy usage of the fault-tolerant hardware, considering all modules, quantum and decoding operations, and algorithmic steps, which to our knowledge is unique to RRE.
@@
    We identify three key power-consuming units that dissipate energy during fault-tolerant computations. These units include decoding cores with a fixed TDP of $P^{\text{decoding}}_{\text{per-core}}=100\text{W}$ on duty during all QEC operations. We assume a fixed number of concurrent decoding cores equal to ${\rm cores}_{{\rm consump}}^{{\rm decoding}}$ are assigned to all decoding opertions, which includes graph preparation, consumption, handover, and T-distillation across full $t^{\text{FT}}_{\text{total}}$. This means we can approximate an upper bound to decoding energy consumption per core as $P^{\text{decoding}}_{\text{per-core}}t^{\text{FT}}_{\text{total}}$.
@@
    We also include 4K units with a total TDP of $P^{\text{tot}}_{\text{4K}}$ and an assumed cooling efficiency factor of $\eta_{\text{4K}}=500$. Additionally, we have lower-temperature MXC units with a total TDP of $P^{\text{tot}}_{\text{MXC}}$ and an assumed cooling efficiency factor of $\eta_{\text{MXC}}=10^{9}$. Each unit consists of several power-consuming stages, and we determine these values through thermal analysis of the signal chain. Overall, we obtain $E_{tot}=(P^{\text{decoding}}_{\text{per-core}}\text{cores}_{\text{distill}}^{\text{decoding}}+\eta_{\text{4K}}P^{\text{tot}}_{\text{4K}}+\eta_{\text{MXC}}P^{\text{tot}}_{\text{MXC}})t^{\text{FT}}_{\text{total}}$.
A.2 Widgetization via Subcircuit Dependency Graph
@@
As described in I.4.2, our resource estimation tooling supports widgetization of Cirq circuits. The primary reason for such a widgetization tool is that fully unrolling a large circuit to its single and two-qubit operations, for a utility-scale application, can place severe pressure on memory and compute resources. Moreover, for resource estimation using RRE, it is unnecessary and inefficient to revisit, re-examine, and recompile subcircuits that have already been prepared for consumption by RRE.
To understand the widgetization process, a few key concepts about the nature of Cirq circuits need to be defined.
A.2.1 Structure of Cirq Circuits
@@
For the purposes of this discussion, we will follow the definitions of quantum circuit objects provided by the documentation for Cirq. As such, we define a quantum gate as “an effect that can be applied to a collection of qubits.” A quantum operation is defined to be a pair consisting of a quantum gate, and a list of qubits that the gate acts on. A moment is a collection of operations, each acting on a disjoint set of qubits. If one imagines a time-ordered list of operations, a moment corresponds to a single time “slice” where operations are being carried out in parallel. Finally, a quantum circuit is a time-ordered list of operations that can be organized, if one wishes, into a time-ordered list of moments.
@@
In addition to a library of common gates, Cirq allows users to define custom gates by providing either a matrix specifying the unitary matrix defining the affect of the gate has on qubits, or by providing instructions on how to create an ordered list of operations from a list of qubits to act on. The latter of the two options can be thought of as instructions on how to “decompose” a gate into constituent gates By providing a method to define gates in terms of other gates, Cirq naturally provides the infrastructure to create circuits that are a hierarchy of nested subcircuits. We will focus on circuits comprised of this latter type of “decomposable” gates.
Utilizing this nested subcircuit structure, a user may construct a relatively simple-looking circuit that consists of billions if not trillions of single and two-qubit gates when fully unrolled to an alphabet of indivisible operations. This is particularly true where some subcircuits consist of many repetitions of smaller subcircuits. The strategy executed during widgetization is to find moderately sized, repeating, subcircuits (called “widgets”) by recursively dividing and analyzing this nested subcircuit hierarchy.
An example of a nested circuit structure is in fig. 16.
@@
One may define a measure of subcircuit size, such as the number of gates or the number of active qubits. Using this measure, one may decide that any particular subcircuit is “too large” to be a “widget”, and therefore the subcircuit needs to be divided into smaller constituent subcircuits. We find it helpful to define two different methods for dividing a subcircuit into its constituents.
The first, and most natural, is to “decompose” the subcircuit into operations by using the decomposition instructions provided by the subcircuit’s gates. This allows the subcircuit to be divided into constituent subcircuits, each comprised of a single operation. An example of this type of splitting is shown in fig. 16(a) and fig. 16(c).
The second type of division is to “slice” the subcircuit into collections of moments. The second method becomes essential when a subcircuit that involves many active qubits decomposes into many (thousands, tens of thousands) of operations, each acting on a few qubits. For example, a circuit consisting of thousands of Toffoli gates, each acting on 3 of a hundred qubits. In these cases the sheer number of constituent operations slows analysis down, and provides subcircuits that are far too small to be considered widgets. By slicing, the subcircuit can be divided into fewer constituent subcircuits of moderate size. An example of this type of division is shown in fig. 16(b).
@@
To accelerate the widgetization process, the user can provide rules that map subcircuits to unique identifiers defining classes of “graph state equivalent” subcircuits. If two subcircuits are mapped to the same identifier, it is assumed that they will be compiled to graph states consuming equivalent resources. Only one graph state equivalence class representative must be split further and/or compiled. This offers valuable savings in the amount of classical processing time when performing both widgetization and resource estimation.
(a) Example algorithm consisting of three subcircuits; Subcircuit A (magenta) that is repeated 4 times and Subcircuit B (yellow) that occurs before and after the 4 repetitions of Subcircuit B.
(b) Splitting of Subcircuit B into three smaller subcircuits via slicing. In this example, these slices meet the criterion not to require further splitting and, as such, are considered to be widgets.
@@
A.2.2 Widgetization
To widgetize without fully unrolling down to single and two qubit operations, we perform the following steps:
@@ A.2.2 Widgetization
    Construction of a subcircuit dependency graph. This describes the nested structure of the circuit as a directed graph.
@@
A.2.3 Construction of the Subcircuit Dependency Graph
Figure 17: Subcircuit dependency graph of the example circuit described in fig. 16. Dashed lines indicate subgraphs that are graph state equivalent to a previously seen subcircuit and, therefore, are not directly stored in the dependency graph. Quantities in brackets are active qubits in the subcircuit. Numbers in square brackets indicate the ordering of the edges, while numbers in angle brackets indicate repetitions. An example of a splitting criterion that may have created this graph is to split a subcircuit if its number of active qubits is greater than or equal to 400.
The first step in the widgetization process is constructing a directed graph (not to be confused with algorithm-specific graph states) that represents the original circuit’s nested subcircuit structure. In this directed graph, the vertices represent subcircuits. A directed edge connects a parent vertex to a child vertex if the child vertex represents one of its parent’s constituent subcircuits. Directed edges between parent and child vertices are ordered to execute the subcircuits. This graph is created by recursively splitting subcircuits into smaller ones. This recursive splitting continues until a subcircuit meets specific user-defined criteria. The vertices that represent subcircuits and are not split are called leaf vertices.
An example of a criterion that would prevent a subcircuit from being split is if its number of qubits or $T$-gates falls below a user-specified threshold. This recursive splitting is carried out in a depth-first search manner.
Figure 17 shows an example of a graph created from the example circuit described in fig. 16.
A.2.4 Enumeration of Widgets and Stitches
Once a directed graph representing the nested structure of the original circuit has been created, it is then used to identify widgets and enumerate how many times they would appear during circuit execution. To account for teleportation overheads, we also count the number of times any ordered pair of widgets (called “stitches”) would be required during execution.
@@ A.2.4 Enumeration of Widgets and Stitches
From the standpoint of formal language theory, circuits constructed this way can be seen as variables in a Context-Free Grammar, with the productions representing the gate decomposition rules. At the same time, the Subcircuit Dependency Graph can be regarded as a parse tree. Widgets are viewed as terminals, and stitches are interpreted as ordered pairs of terminals. The number of widgets and stitches is then counted based on their occurrence in the yield of the parse tree.
A.3 A unitariness verification protocol for RRE outputs
The robust verification of our compilation and estimation methodology at different layers was a priority when designing RRE. To this end, we have developed an exact verification protocol and a series of tests to precisely check the correctness of individual subgraphs, $\{g\}$, and connected schedules, $\{S_{\text{consump}}\}$ and $\{S_{\text{meas}}\}$, as long as the widgets originated from small circuits with a few logical qubits. The protocol involves a step for exact simulation and matrix manipulation for the quantum algorithm, leading to the widget size restriction. Recall that every logical qubit involving nontrivial non-Clifford operations can correspond to 10’s of physical qubits that need exact simulations (one can extend our protocol to 10’s of logical qubits by performing near-exact tensor-network-based simulations).
@@ A.3 A unitariness verification protocol for RRE outputs
$\begin{split}&\{g_{i},S_{\text{frames}}^{g_{i}},S_{\text{input}}^{g_{i}},S_{\text{output}}^{g_{i}},n^{g_{i}}_{\text{logical}},S_{\text{consump}}^{g_{i}},S_{\text{meas}}^{g_{i}},\\ &\ket{\text{state}}^{\text{Cabaliser}}_{i}\}~{}\text{for}~{}i\in\{0,\dots,n_{\text{widgets}}-1\}~{}.\end{split}$ (30)
Recall that RRE, by default, initializes from an all-zero state as |$\ket{\text{input}}\equiv\ket{s_{0}=0,s_{1}=0,\cdots,s_{n_{\text{input}}-1}=0}$. Moreover, the outputted estimation results by RRE for $U$ do not play a role in this exact verification protocol. Now, entirely independent of the operations of the compiler, one can construct the exact inverse of the original unitary $U$ for small-size cases through Kronecker matrix multiplications, namely $U^{\dagger}$. (As an example, if the widget $U_{1}$ is the unitary for a logical QFT4, we then need to construct $U_{1}^{\dagger}$, i.e. InverseQFT4, manually.) If we now apply $U^{\dagger}$ to the final output state that RRE was generated through Cabaliser, |$\ket{\text{output}}\equiv\ket{\text{state}}^{\text{Cabaliser}}_{n_{\text{widget}}-1}$, by unitary principle we should get back the initial state. That is $U^{\dagger}\ket{\text{state}}^{\text{Cabaliser}}_{n_{\text{widgets}}-1}=\ket{s_{0}=0,s_{1}=0,\cdots,s_{n_{\text{input}}-1}=0}$ will verify the correctness of graphs and schedules generated by RRE.
In the RRE test modules and procedures, we simulated and verified the graphs and schedules generated by the software, following the aforementioned protocol for small input algorithms. This encompasses widgets featuring non-Clifford instructions like QFT3 and TOFFOLI3, in addition to various Clifford operations.

### v3

@@ Appendix A Glossary of default RRE outputs
    $t^{\text{decode-delay}}_{tot}$ represents the total decoding delay accumulated from all graph consumption steps, which we must add to the overall FT hardware time. We assume there are no decoying delays during subgraph hand-over operations due to $t_{\text{tock}}^{\text{decoding}}\leq 8t_{\text{inter}}d$. We can determine the total decoding delay, given that $t_{\text{tock}}^{\text{decoding}}>t_{\text{tock}}^{\text{quantum}}$ and $t_{\text{tock}}^{\text{decoding}}>8Ct$, using $t^{\text{decode-delay}}_{\text{tot}}=\left(8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{prep}}\right)\left(t_{\text{tocks}}^{\text{decoding}}-t_{\text{tock}}^{\text{quantum}}\right)+\left\lceil\frac{t^{\text{delay}}_{\text{distill}}}{8Ct}\right\rceil\left(t_{\text{tocks}}^{\text{decoding}}-8Ct\right)$.
@@
  • Parameter 47: Total wall-time over a single FT algorithm step, $t_{\text{hardware}}^{\text{total}}$
@@
    $t_{\text{hardware}}^{\text{total}}$ is our estimated upper-bound for the wall-time accumulated from all non-simultaneous components running in modules on both legs and coherent interconnects of the architecture of section˜II.3 over a single FT algorithm step, $U$. Algorithmic step should not be confused with $n_{\text{widgets}}$ widgetization steps of $U$. (An algorithmic step occurs when the user requires $n_{\text{algo-reps}}$ repetitions of $U$ on the FTQC.) This wall time includes all inter- and intra-tocks required at the consumption stages, delays for graph preparation and $T$-distillation, subgraph handover time, and decoding delay. In other words, this is our overall time cost for one step, which can be formulated as $t_{\text{hardware}}^{\text{total}}=t^{\text{total}}_{\text{intra}}+t^{\text{total}}_{\text{handover-inter}}+t_{\text{decode-delay}}^{\text{total}}$. See section˜II.4.7 for details.
  • Parameter 48: Total FT algorithm time, $t_{\text{FT}}^{\text{total}}$
@@
    RRE allows users to specify the number of times, $n_{\text{algo-reps}}$, with which they would like to repeat an exact copy of the complete algorithm $U$ back-to-back. For these cases, the total FT time is calculated as $t_{\text{FT}}^{\text{total}}=n_{\text{algo-reps}}t_{\text{hardware}}^{\text{total}}$.
  • Parameter 49: Total energy consumption, $E_{\text{tot}}$
@@
    We report estimates for $E_{\text{tot}}$, the upper-bound on the total energy usage of the FTQC, considering all modules, quantum and decoding operations, and algorithmic steps, which, to our knowledge, is unique to RRE.
@@
    We identify three key power-consuming units that dissipate energy during FT computations. These units include decoding cores with a fixed TDP of $P^{\text{decoding}}_{\text{per-core}}=100\text{W}$ on duty during all QEC operations. We assume a fixed number of concurrent decoding cores equal to ${\rm cores}_{{\rm consump}}^{{\rm decoding}}$ are assigned to all decoding opertions, which includes graph preparation, consumption, handover, and $T$-distillation across full $t^{\text{FT}}_{\text{total}}$. This means we can approximate an upper bound to decoding energy consumption per core as $P^{\text{decoding}}_{\text{per-core}}t^{\text{FT}}_{\text{total}}$.
@@
    We also include 4 K units with a TDP of $P_{\text{dissip},\text{4K}}$ and assume a single cooling efficiency factor of $\eta_{\text{4K}}=500$ for this stage. Additionally, we have lower-temperature MXC (20 mK) units with a TDP of $P_{\text{dissip},\text{20mK}}$ and assume a single cooling efficiency factor of $\eta_{\text{20mK}}=10^{9}$ for this stage. Each unit comprises several power-consuming stages, and we determine these values via thermal analysis of the signal chain (see section˜II.3.3 and [77, 48, 67] for details). Overall, we obtain $E_{tot}=(P^{\text{decoding}}_{\text{per-core}}\text{cores}_{\text{distill}}^{\text{decoding}}+\eta_{\text{4K}}P_{\text{dissip},\text{4K}}$ $\eta_{\text{20mK}}P_{\text{dissip},\text{20mK}})t^{\text{FT}}_{\text{total}}$.
Appendix B Widgetization via Subcircuit Dependency Graph
@@
In section˜II.4.3, we established that our resource estimation tooling supports the widgetization of Cirq circuits. The primary reason for such a widgetization tool is that fully unrolling a large circuit into its single- and two-qubit operations for a utility-scale application can place severe pressure on memory and compute resources. Moreover, for resource estimation using RRE, it is unnecessary and inefficient to revisit, re-examine, and recompile subcircuits that have already been prepared for RRE’s use. To understand the widgetization process, a few key concepts about the nature of Cirq circuits need to be defined.
B.0.1 Structure of Cirq circuits
@@
For this discussion, we will follow the definitions of quantum circuit objects in the Cirq documentation. As such, we define a quantum gate as “an effect that can be applied to a collection of qubits” (logical qubits for our purpose). A quantum operation is determined to be a pair consisting of a quantum gate and a list of logical qubits that the gate acts on. A moment is a collection of operations, each acting on a disjoint set of logical qubits. If one imagines a time-ordered list of operations, a moment corresponds to a single time “slice” where operations are being carried out in parallel. Finally, a quantum circuit is a time-ordered sequence of operations that can, if desired, be organized into a time-ordered sequence of moments.
@@
In addition to a library of common gates, Cirq allows users to define custom gates by providing either a matrix specifying the unitary matrix defining the effect of the gate on logical qubits, or by providing instructions on how to create an ordered list of operations from a list of qubits to act on. The latter of the two options can be thought of as instructions for “decomposing” a gate into constituent gates. By providing a method to define gates in terms of other gates, Cirq naturally provides the infrastructure to create circuits that are a hierarchy of nested subcircuits. We will focus on circuits comprised of this latter type of “decomposable” gates.
Utilizing this nested subcircuit structure, a user may construct a relatively simple-looking circuit that consists of billions, if not trillions, of single and two-qubit gates when fully unrolled to an alphabet of indivisible operations. This is particularly true where some subcircuits consist of many repetitions of smaller subcircuits. The strategy employed during widgetization is to identify moderately sized, recurring subcircuits (called “widgets”) by recursively decomposing and analyzing the nested subcircuit hierarchy. An example of a nested circuit structure is in fig.˜14.
@@
One may define a measure of subcircuit size, such as the number of gates or active logical qubits. Using this measure, one may decide that any particular subcircuit is “too large” to be a “widget”, and therefore the subcircuit needs to be divided into smaller constituent subcircuits. We find it helpful to define two distinct methods for dividing a subcircuit into its constituent parts.
The first, and most natural, is to “decompose” the subcircuit into operations by using the decomposition instructions provided by the subcircuit’s gates. This allows the subcircuit to be divided into constituent subcircuits, each comprised of a single operation. An example of this type of splitting is shown in fig.˜14(a) and fig.˜14(c).
The second type of division is to “slice” the subcircuit into collections of moments. The second method becomes essential when a subcircuit that involves many active logical qubits decomposes into many (thousands or tens of thousands) of operations, each acting on a few qubits. For example, a circuit consisting of thousands of Toffoli gates, each acting on 3 of a hundred logical qubits. In these cases, the sheer number of constituent operations slows analysis down and provides subcircuits that are far too small to be considered widgets. By slicing, the subcircuit can be divided into fewer constituent subcircuits of moderate size. An example of this type of division is shown in fig.˜14(b).
@@
(a) Example algorithm consisting of three subcircuits. Subcircuit A (magenta) that is repeated 44 times and Subcircuit B (yellow) that occurs before and after the 44 repetitions of Subcircuit B.
(b) Splitting of Subcircuit B into three smaller subcircuits via slicing. In this example, these slices meet the criterion for not requiring further splitting and, as such, are considered widgets.
@@
To accelerate the widgetization process, the user can provide rules that map subcircuits to unique identifiers, defining classes of “graph-state-equivalent” subcircuits. If two subcircuits are mapped to the same identifier, it is assumed that they will be compiled into graph states that consume equivalent resources. Only one representative of each graph state equivalence class must be split further and/or compiled. This yields substantial savings in classical processing time when performing both widgetization and resource estimation.
B.0.2 Widgetization
To widgetize $U$ without fully unrolling down to single and two-qubit operations, we perform the following steps:
@@ B.0.2 Widgetization
    Construction of a subcircuit dependency graph. This describes the circuit’s nested structure as a directed graph.
@@
B.0.3 Construction of the Subcircuit Dependency Graph
Figure 15: Subcircuit Dependency Graph of the example circuit described in fig.˜14. Dashed lines indicate subgraphs that are graph state equivalent to a previously seen subcircuit and, therefore, are not directly stored in the dependency graph. Quantities in brackets are active qubits in the subcircuit. Numbers in square brackets indicate the ordering of the edges, while numbers in angle brackets indicate repetitions. An example of a splitting criterion that may have created this graph is to split a subcircuit if its number of active qubits is greater than or equal to $400$.
The first step in the widgetization process is constructing a directed graph (not to be confused with ASG states) that represents the original circuit’s nested subcircuit structure. In this directed graph, the vertices represent subcircuits. A directed edge connects a parent vertex to a child vertex if the child vertex represents one of its parent’s constituent subcircuits. Directed edges between parent and child vertices are ordered to execute the subcircuits. This graph is created by recursively splitting subcircuits into smaller ones. This recursive splitting continues until a subcircuit meets specific user-defined criteria. The vertices that represent subcircuits and are not split are called leaf vertices.
An example of a criterion that would prevent a subcircuit from being split is if its number of logical qubits or $T$-gates falls below a user-specified threshold. This recursive splitting is performed in a depth-first manner.Figure 15 shows an example of a graph created from the example circuit described in fig.˜14.
B.0.4 Enumeration of widgets and stitches
Once a directed graph representing the nested structure of the original circuit is created, it is used to identify widgets and count how many times they appear during circuit execution. To account for teleportation overheads, we also count the number of times any ordered pair of widgets (called “stitches”) would be required during execution.
@@ B.0.4 Enumeration of widgets and stitches
From the standpoint of formal language theory, circuits constructed this way can be viewed as variables in a Context-Free Grammar, with the productions corresponding to the gate decomposition rules. At the same time, the Subcircuit Dependency Graph can be regarded as a parse tree. Widgets are viewed as terminals, and stitches are interpreted as ordered pairs of terminals. The number of widgets and stitches is then counted based on their occurrence in the yield of the parse tree.
Appendix C A unitariness verification protocol for RRE outputs
Robust verification of our compilation and estimation methodology across multiple layers was a priority in the design of RRE. To this end, we have developed an exact verification protocol and a series of tests to precisely check the correctness of individual subgraphs, $\{g\}$, and connected schedules, $\{S_{\text{consump}}\}$ and $\{S_{\text{meas}}\}$, as long as the widgets originated from small circuits with a few logical qubits. The protocol includes a step for exact simulation and matrix manipulation in the quantum algorithm, which imposes a widget-size restriction. Recall that every logical qubit involving nontrivial non-Clifford operations can correspond to 10’s of physical qubits that need exact simulations (one may extend our protocol to 10’s of logical qubits by performing near-exact tensor-network-based simulations).
@@ Appendix C A unitariness verification protocol for RRE outputs
[…]
In the RRE test modules and procedures, we simulated and verified the graphs and schedules generated by the software, following the protocol above for selected small input algorithms. These tests included widgets with non-Clifford instructions such as QFT3 and TOFFOLI3, in addition to various Clifford operations.
Appendix D Noise modeling and logical error rate scaling
Section II.3.1 established that our resource estimation framework requires a physical-to-logical error scaling law to obtain resource counts. This error scaling law is given by, eq.˜1, $p_{C}=\kappa\left(\frac{p}{p_{\text{thresh}}}\right)^{(d+1)/2},$ where $p_{C}$ is the logical error rate per cycle, $p$ is the homogeneous physical error rate, and $\kappa$ and $p_{\text{thresh}}$ are the coefficients to be fitted. Below, we describe the computational pipeline we used to determine the scaling coefficients, as well as the exemplar device noise model used for the results in this work.
D.1 Device noise model
For a preliminary demonstration of RRE, we chose a simple noise model that also shows the richness and flexibility the tool offers researchers when considering other devices, circuits, and algorithms. To this end, we split the noise model into two classes: one simple model for single-qubit gates and another, more realistic model for two-qubit gates.
For single-qubit gates, we chose a depolarizing noise channel, parameterized by the homogeneous physical error rate $p$. Admittedly, this choice sacrifices realism for ease of implementation, but for an initial baseline demonstration, we considered it appropriate. The noise model for two-qubit gates was chosen to be weight-2 Pauli-error channels, with error rates derived from device-specific gate simulations. As mentioned in section˜II.3.1, this manuscript obtains resource counts for FT execution of circuits protected by a square planar $iSWAP$ surface code. As iSWAP gates were the only two-qubit gate in the surface code, the Pauli Error channel rates were derived from master equation simulations of parametric $iSWAP$ gates. The simulations were based on the two-qubit system with one tunable coupler described in [81]. Process tomography was carried out to obtain the Pauli Error rates from the difference between the ideal $iSWAP$ Pauli Transfer Matrix and the simulated Pauli Transfer Matrix.
D.2 Scaling coefficient pipeline
Figure 16: The Logical Error Rate (LER) to Physical Error Rate (PER) scaling obtained via the Scaling Coefficient pipeline in Appendix D D.2. Dots are values calculated by the pipeline, and solid lines are lines of best fit, resulting in the power law given in eq.˜1 with fit parameters of $\kappa[\text{MWPM}]=0.009$ and $p_{\text{thresh}}[\text{MWPM}]=0.016$.
To determine the coefficients for the logical-to-physical error rate scaling law, eq.˜1, we performed a series of surface code simulations and coefficient fittings by using Sinter [24] and Stim [22].
Given the stochastic nature of the simulations, we conducted convergence studies to ensure the stability of the coefficients across a large number of shots. This was done by simulating and fitting coefficients for an increasing number of shots and a maximum number of errors, until the coefficients converged. The number of shots and maximum number of errors were $2^{k}\times 100,000$ and $2^{k}\times 10000$ respectively, where $k$ ranged from $1$ to $11$.
For each fixed number of shots and maximum number of errors, we swept over 20 values for $p$ ranging from $10^{-2}$ to $10^{-4}$ and distances of $d\in\{3,5,7,9\}$. For each $p$ and distance combination, noisy surface code circuits were simulated for $10\times d$ rounds (cycles). Noisy circuits were constructed by starting with a noiseless $iSWAP$ surface code circuit written in Stim provided by [23]. We then used the aforementioned noise model to add noise to this circuit in Stim. Syndromes were decoded using pymatching MWPM decoder [29], followed by the calculations of the logical error rates (per cycle). Observed data were then fit to the two-coefficient scaling law. Fitted coefficients were observed to converge to two significant figures by $k=11$, and the coefficients for $k=11$ were used in all downstream resource estimates of section˜IV. The fit and logical error rates for $k=11$ are shown in fig.˜16.